OpenStudy (anonymous):

PLEASE JUST CHECK ! D ! MEDAL MEDAL MEDAL Kevin wants to make a mixture that is 50% lemon juice and 50% lime juice. How much 100% lemon juice should he add to a juice that is 40% lemon juice and 60% lime juice to make 6 gallons of the 50% lemon/50% lime juice mixture? 1 gallon 0.5 gallon 2 gallons 2.5 gallons

OpenStudy (anonymous):

ANSWER IS A ! A ! A ! A ! A !

OpenStudy (anonymous):

A is my corrected answer sorry i put d thats why...

OpenStudy (anonymous):

AAAAAAAAAAAAAA

OpenStudy (anonymous):

omg thaanks again !!!!!!!!!!! @TaylorSimonelli

OpenStudy (anonymous):

Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? Write a system of equations to model the problem if x represents Joelle’s score, and y represents Anita’s score. x + y = 425 y = 2x – 40 x + y = 425 y = 2x + 40 x – y = 425 y = 2x – 40 x + y = 40 y = 2x – 425 ANSWER A !!! WOULD THIS BE CORRECT TO ? ? ? @TaylorSimonelli